OpenStudy (anonymous):

Find the value of x and y where x and y are real numbers. 4+(x+2y)i=x+2i

OpenStudy (anonymous):

wow...complex nos. wat do u want?

OpenStudy (anonymous):

ohh..edited..!!!

OpenStudy (anonymous):

:)

OpenStudy (jessiegonzales):

is there a multiple choice?

OpenStudy (anonymous):

no but i know the answer

OpenStudy (jessiegonzales):

oh then why did u post this? Whats the answer?

OpenStudy (anonymous):

x=4, y=-1 & i posted it because I don't know how to do it..

OpenStudy (anonymous):

@ganeshie8 please help

OpenStudy (anonymous):

Anybody..

OpenStudy (anonymous):

u can compare d real n imaginary part...!!!

OpenStudy (anonymous):

@Arababeee hey...wat do u think?

OpenStudy (anonymous):

\[4+(x+2y)i=x+2i\] In this equation you have one complex number equal to another complex number. They must be the same complex number, right? For this to be true, both the real and imaginary parts of both sides must be the same. That is, for two complex numbers \(a+bi\) and \(c+di\), the numbers are equal if and only if \(a=c\) and \(b=d\). In this case, setting the real parts equal gives us \[4=x\] and setting the imaginary parts equal gives \[x+2y=2\] Solve for \(x\) and \(y\). You already know the value of \(x\), so substitute into the second equation and find \(y\).